[figure: 9.7 cm]

EFFECT OF NEUTRON IRRADIATION ON THE EPR SPECTRA OF CuGeO_{3}

E. Mosiniewicz-Szablewska^{a}, L. Gladczuk^{a}, H.A. Dabkowska^{b}, B. Pytel^{c}

and H. Szymczak^{a} aInstitute of Physics, Polish Academy of Sciences, Warsaw, Poland

bBrockhouse Institute for Materials Research, McMaster University, Hamilton, Canada cInstitute of Atomic Energy, Swierk, Poland

The influence of neutron irradiation on the electron paramagnetic resonance (EPR) spectra of the spin-Peierls compound CuGeO_{3} has been investigated in the wide temperature range 2 - 300 K. It was found that the irradiation induced appreciable changes in the EPR signal intensity, resonance linewidth and g-factor of this material. It was also shown that the irradiation leads to a decrease of the spin-Peierls transition temperature. For higher irradiation doses, an antiferromagnetically ordered phase can be observed at lower temperatures. This phase coexists with the spin-Peierls phase. The observed decrease in the spin-Peierls transition temperature after irradiation may be associated with some changes in topological and chemical short-range order or with the transmutation of the fraction of ^{63}Cu in the ^{63}Cu(n,α)^{60}Co reaction.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Ewa Mosiniewicz-Szablewska

Address for correspondence : Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : mosin@ifpan.edu.pl